OpenStudy (danieke):

fan and medal What is the equation of a line passing through (-8, -3) and having a slope of -4?

OpenStudy (danieke):

@cookielate

OpenStudy (cookielate):

Do you know how to find it or not ?

OpenStudy (danieke):

not

OpenStudy (anonymous):

Do you got this cookie?

OpenStudy (cookielate):

you could go ahead and explain it

OpenStudy (danieke):

y = −4x + 20 y = 4x + 20 y = −4x – 35 y = 4x – 35 y = -4x + 4 the options

OpenStudy (cookielate):

The first one because its -4 and the rest have a different equation that is not the solution